A baseball diamond is a square playing surface, meaning the lines (base paths) are at right angles with
respect to one another. Bases are 27.43 meters (90 feet) apart. Throwing across the diamond, as a
catcher does when trying to throw from home plate to second base, is the longest throw from one base to
another. How long is this distance? ►_____________ meters
: A baseball diamond is a square playing surface, meaning the lines (base paths) are at right angles with
respect to one another. Bases are 27.43 meters (90 feet) apart. Throwing across the diamond, as a
catcher does when trying to throw from home plate to second base, is the longest throw from one base to
another. How long is this distance? ►_____________ meters

Answer by stanbon(18991)

Bases are 27.43 meters
----------------------------
Let the diagonal be "x" meters.
Use Pythagoras to solve for "x".
-------------------------
x^2 = 27.43^2 + 27.43^2
x^2 = 2*27.43^2
x = 27.43sqrt(2) meters
==========================
Cheers,
Stan H.